OpenStudy (anonymous):

What is the mean of the following set? 1/2 1/8 1/4 5/8 7/12 1/4 will give a medal to who ever answers correctly

OpenStudy (cwrw238):

add up all the fractions and divide by 6

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

A handy tip from someone long ago: \[\frac{1}{2} + \frac{1}{8} + \frac{1}{4} + \frac{5}8 + \frac{7}{12} + \frac{1}4 = x\]Multiply by the least common multiple of 2,4,8,12, which is 24: \[\frac{24}{2} + \frac{24}{8} + \frac{24}4 + \frac{120}8 + \frac{168}{12} + \frac{24}{4} = 24x\]\[12+3+6+15+14+6=24x\]\[56=24x\]\[x=\]